# An Efficient Local Search Approach for Polarized Community Discovery in Signed Networks

### v1

@@ B.3 Complete Results on All Datasets
(a) Epinions, $k=2$, $\beta=0.1$
(b) Wikipol, $k=6$, $\alpha=\frac{1}{k-1}$
Figure 2: Investigation of the impact of $\alpha$ and $\beta$.
In Figure 2, we illustrate the impact of varying $\alpha$ and $\beta$. According to Figure 2(a), increasing $\alpha$ naturally balances intra-cluster cohesion and inter-cluster opposition. The size proportion, defined as the fraction of non-neutral objects in $V$, remains constant as $\alpha$ varies. Figure 2(b) shows that increasing $\beta$ monotonically reduces the number of non-neutral objects, leading to denser clusters, as indicated by improved MAC and MAO scores. Notably, balance remains stable across different $\beta$ values, unlike the baseline SCG-MA.
Table 4: Results for different methods and datasets w.r.t. the balance-aware polarity metric. $|E|$ is the number of non-zero edges and $|E_{-}|$ is the number of negative edges. We observe that our method (LSPCD) often yields the highest scores.
W8 BTC WikiV REF SD WikiC EP WikiP
$|V|$ 790 5.9K 7.1K 10.9K 82.1K 116.7K 131.6K 138.6K
$|E|$ 116K 214.5K 1M 251.4K 500.5K 2M 711.2K 715.9K
$|E_{-}|/|E|$ 0.2 0.2 0.2 0.1 0.2 0.6 0.2 0.1
$k=2$ LSPCD (ours) 19.4 10.9 13.9 61.8 9.7 102.2 56.2 12.4
SCG-MA 0.4 2.4 0.4 1.3 0.8 44.7 2.5 0.6
SCG-MO 0.5 0.6 0.4 1.3 0.7 39.6 2.4 0.4
SCG-B 4.9 19.7 0.7 1.7 1.5 47.3 3.8 0.9
SCG-R 3.4 1.8 4.6 2.1 1.1 31.9 6.6 3.0
KOCG-top-$1$ 10.4 1.0 3.8 4.8 1.3 3.5 3.3 1.8
KOCG-top-$r$ 11.3 3.3 2.2 11.7 2.2 2.9 10.1 1.1
BNC-$(k+1)$0.4 -0.9 -0.6 -1.0
BNC-$k$ 0.7 0.0 0.0 0.0
SPONGE-$(k+1)$ 19.7 0.6 0.3 0.6
SPONGE-$k$ 22.4 0.0 0.0 0.0
$k=4$ LSPCD (ours) 6.9 3.6 5.9 3.9 6.7 26.3 14.9 5.4
SCG-MA 5.8 0.3 0.8 0.3 1.9 0.7 0.1 0.7
SCG-MO 4.9 0.3 0.4 0.4 0.2 6.6 0.1 0.0
SCG-B 0.3 0.1 0.2 0.1 0.5 18.5 0.2 0.1
SCG-R 1.3 0.9 0.9 0.6 2.4 2.0 0.3 1.0
KOCG-top-$1$ 5.8 4.2 1.2 2.6 0.9 0.4 4.0 0.9
KOCG-top-$r$ 5.4 2.7 2.6 1.6 1.0 2.8 5.9 1.2
BNC-$(k+1)$0.1 -0.8 -0.4 -1.0
BNC-$k$ 0.5 0.0 0.0 0.0
SPONGE-$(k+1)$ 22.9 0.0 0.4 0.3
SPONGE-$k$ 18.6 0.0 0.0 0.0
$k=6$ LSPCD (ours) 4.4 2.2 3.6 2.2 3.8 16.4 6.8 3.5
SCG-MA 3.1 0.1 0.2 0.1 0.0 0.6 0.4 0.0
SCG-MO 3.6 0.1 0.2 0.1 0.1 4.2 0.1 0.0
SCG-B 0.3 0.2 0.2 0.1 0.2 1.8 0.2 0.1
SCG-R 0.3 0.3 0.4 0.2 0.4 2.3 0.1 0.0
KOCG-top-$1$ 6.1 1.7 2.5 3.3 1.7 0.4 2.7 3.7
KOCG-top-$r$ 5.0 1.2 1.8 2.5 1.3 1.1 2.8 0.7
BNC-$(k+1)$0.1 -0.1 -0.8 -0.4
BNC-$k$ 0.5 0.0 0.0 0.0
SPONGE-$(k+1)$ 17.7 0.0 0.4 0.6
SPONGE-$k$ 16.0 0.0 0.0 0.0
Table 5: Results for different methods and datasets w.r.t. the polarity metric. $|E|$ is the number of non-zero edges and $|E_{-}|$ is the number of negative edges. Our method (LSPCD) yields the best or competitive results, despite the fact that it, unlike SCG methods, does not directly optimize polarity.
W8 BTC WikiV REF SD WikiC EP WikiP
$|V|$ 790 5.9K 7.1K 10.9K 82.1K 116.7K 131.6K 138.6K
$|E|$ 116K 214.5K 1M 251.4K 500.5K 2M 711.2K 715.9K
$|E_{-}|/|E|$ 0.2 0.2 0.2 0.1 0.2 0.6 0.2 0.1
$k=2$ LSPCD (ours) 223.4 29.0 62.3 146.1 75.9 190.8 127.8 82.0
SCG-MA 236.6 28.8 71.5 172.2 77.5 155.2 128.3 82.8
SCG-MO 236.6 29.5 71.7 174.1 79.7 175.7 128.7 88.4
SCG-B 200.6 21.6 37.6 116.3 61.0 129.3 156.4 46.5
SCG-R 214.6 14.2 54.7 120.9 29.7 101.1 72.3 36.1
KOCG-top-$1$ 13.0 1.0 7.6 11.6 2.0 5.9 8.2 3.0
KOCG-top-$r$ 13.0 3.8 2.3 15.4 2.6 3.4 14.0 1.3
BNC-$(k+1)$0.7 -10.8 -1.1 -1.0
BNC-$k$ 184.6 5.3 15.8 41.5
SPONGE-$(k+1)$ 88.0 1.0 1.0 1.0
SPONGE-$k$ 191.4 5.1 15.8 41.5
$k=4$ LSPCD (ours) 218.5 23.3 52.6 139.2 61.1 113.6 111.5 71.6
SCG-MA 205.1 25.1 52.9 94.5 35.5 104.9 127.4 56.5
SCG-MO 213.2 25.3 53.1 82.1 38.5 117.9 129.0 39.7
SCG-B 211.6 12.4 24.8 116.2 48.3 49.8 94.4 45.7
SCG-R 214.6 8.0 19.5 118.7 10.7 41.1 65.1 33.7
KOCG-top-$1$ 9.1 8.4 4.5 15.0 2.6 4.5 8.9 3.1
KOCG-top-$r$ 7.4 5.0 3.3 3.7 3.0 3.8 11.0 4.4
BNC-$(k+1)$0.2 -9.4 -1.1 -1.0
BNC-$k$ 185.3 5.2 15.8 41.5
SPONGE-$(k+1)$ 53.8 1.1 1.0 1.0
SPONGE-$k$ 71.2 5.1 15.8 41.5
$k=6$ LSPCD (ours) 217.3 20.0 46.2 137.6 57.1 96.1 103.4 58.7
SCG-MA 207.3 14.6 45.5 84.9 37.8 102.6 88.8 57.5
SCG-MO 205.8 15.2 47.0 55.6 34.6 111.6 129.2 41.8
SCG-B 211.6 9.3 23.3 116.2 47.7 46.1 94.5 46.0
SCG-R 201.2 6.9 10.4 50.3 7.9 18.3 43.3 3.3
KOCG-top-$1$ 7.9 4.1 4.5 8.6 3.6 4.9 6.0 10.1
KOCG-top-$r$ 9.1 3.6 3.1 4.0 3.3 1.5 6.8 3.6
BNC-$(k+1)$0.2 -4.2 -1.1 -0.8
BNC-$k$ 185.2 5.2 15.8 41.5
SPONGE-$(k+1)$ 47.8 1.3 1.0 1.0
SPONGE-$k$ 57.9 5.1 15.8 41.5
Table 6: Detailed results on the WoW-EP8 dataset from various aspects. LSPCD (avg) and LSPCD (std) respectively indicate the mean and standard deviation across five runs of our method with different seeds.
SIZE BAL K MAC MAO CC+ CC- DENS ISO POL BA-POL RT
$k=2$ LSPCD (avg) 586 0.087 2 0.261 0.098 0.757 0.509 0.511 0.769 223.406 19.409 0.249
LSPCD (std) 0 0.0 0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.034
SCG-MA 527 0.002 1 0.52 0.0 0.762 0.0 0.59 0.725 236.55 0.448 1.025
SCG-MO 517 0.002 1 0.527 0.0 0.769 0.0 0.596 0.708 236.592 0.457 1.026
SCG-B 583 0.025 2 0.274 0.094 0.697 -0.369 0.514 0.767 200.604 4.918 4.657
SCG-R 513 0.016 2 0.272 0.104 0.761 0.451 0.552 0.654 214.616 3.386 0.517
KOCG-top-$1$ 16 0.8 2 0.986 0.889 0.965 0.778 1.0 0.019 13.0 10.4
KOCG-top-$r$ 527 0.869 2 0.467 0.11 0.658 -0.598 0.559 0.691 12.964 11.269
BNC-$(k+1)$ 3 0.667 2 0.5 0.0 -1.0 0.0 0.333 0.007 -0.667 -0.444 0.99
BNC-$k$ 790 0.004 2 0.152 0.049 0.628 1.0 0.372 1.0 184.63 0.702 0.537
SPONGE-$(k+1)$ 375 0.224 2 0.204 0.095 0.738 0.318 0.343 0.346 87.957 19.705 0.581
SPONGE-$k$ 790 0.117 2 0.191 0.093 0.696 0.15 0.372 1.0 191.38 22.404 0.572
$k=4$ LSPCD (avg) 590 0.031 4 0.156 0.082 0.76 0.426 0.506 0.772 218.458 6.862 0.317
LSPCD (std) 1 0.0 0 0.003 0.002 0.001 0.016 0.0 0.002 0.142 0.008 0.044
SCG-MA 599 0.028 4 0.58 0.13 0.762 -0.32 0.527 0.822 205.137 5.828 1.198
SCG-MO 568 0.023 4 0.827 0.141 0.77 -0.349 0.55 0.776 213.211 4.939 1.176
SCG-B 615 0.002 1 0.421 0.0 0.693 0.0 0.498 0.822 211.561 0.343 6.233
SCG-R 503 0.006 4 0.211 0.056 0.762 0.747 0.564 0.644 214.623 1.301 1.131
KOCG-top-$1$ 31 0.636 4 0.962 0.668 0.944 0.339 0.978 0.039 9.054 5.761
KOCG-top-$r$ 599 0.735 4 0.436 0.099 0.692 -0.618 0.516 0.806 7.393 5.434
BNC-$(k+1)$ 8 0.4 4 0.5 0.0 -1.0 0.0 0.036 0.003 -0.25 -0.1 0.62
BNC-$k$ 790 0.003 4 0.327 0.03 0.632 1.0 0.372 1.0 185.305 0.473 0.594
SPONGE-$(k+1)$ 485 0.426 4 0.448 0.06 0.886 -0.53 0.33 0.433 53.823 22.911 0.593
SPONGE-$k$ 790 0.261 4 0.383 0.08 0.803 -0.499 0.372 1.0 71.162 18.585 0.61
$k=6$ LSPCD (avg) 591 0.02 6 0.142 0.071 0.761 0.414 0.505 0.773 217.344 4.419 0.269
LSPCD (std) 0 0.0 0 0.006 0.002 0.0 0.009 0.0 0.001 0.142 0.003 0.015
SCG-MA 598 0.015 6 0.785 0.113 0.763 -0.339 0.527 0.819 207.299 3.141 1.226
SCG-MO 591 0.017 6 0.756 0.171 0.77 -0.321 0.534 0.811 205.796 3.576 1.378
SCG-B 615 0.002 1 0.421 0.0 0.693 0.0 0.498 0.822 211.561 0.343 7.512
SCG-R 744 0.001 5 0.323 0.178 0.669 0.313 0.41 0.978 201.172 0.275 1.295
KOCG-top-$1$ 42 0.778 6 0.992 0.605 0.984 0.303 0.934 0.053 7.905 6.148
KOCG-top-$r$ 598 0.554 6 0.472 0.095 0.73 -0.637 0.522 0.812 9.109 5.045
BNC-$(k+1)$ 10 0.5 6 0.5 0.0 -1.0 0.0 0.022 0.002 -0.2 -0.1 0.977
BNC-$k$ 790 0.003 6 0.552 0.016 0.632 1.0 0.372 1.0 185.198 0.473 0.61
SPONGE-$(k+1)$ 572 0.372 6 0.537 0.065 0.893 -0.511 0.326 0.536 47.762 17.749 0.607
SPONGE-$k$ 790 0.276 6 0.53 0.075 0.834 -0.529 0.372 1.0 57.868 15.99 0.599
Table 7: Detailed results on the Bitcoin dataset from various aspects. LSPCD (avg) and LSPCD (std) respectively indicate the mean and standard deviation across five runs of our method with different seeds.
SIZE BAL K MAC MAO CC+ CC- DENS ISO POL BA-POL RT
$k=2$ LSPCD (avg) 155 0.377 2 0.2 0.143 0.94 0.969 0.199 0.211 29.022 10.947 2.203
LSPCD (std) 0 0.0 0 0.0 0.0 0.003 0.001 0.001 0.001 0.013 0.005 0.161
SCG-MA 179 0.084 2 0.298 0.068 0.906 0.873 0.179 0.216 28.838 2.418 0.122
SCG-MO 138 0.022 2 0.114 0.147 0.91 0.778 0.237 0.184 29.522 0.646 0.123
SCG-B 40 0.909 2 0.248 0.87 0.956 1.0 0.56 0.201 21.65 19.682 2.243
SCG-R 842 0.124 2 0.022 0.009 0.908 0.812 0.019 0.394 14.24 1.763 1.024
KOCG-top-$1$ 2 1.0 2 1.0 1.0 0.0 1.0 1.0 0.004 1.0 1.0
KOCG-top-$r$ 179 0.885 2 0.063 0.055 0.266 0.182 0.094 0.165 3.754 3.324
BNC-$(k+1)$ 50 0.083 2 0.058 0.0 -0.516 0.0 0.425 0.581 -10.76 -0.897 0.341
BNC-$k$ 5881 0.008 2 0.059 0.001 0.721 0.694 0.001 1.0 5.268 0.043 0.202
SPONGE-$(k+1)$ 6 0.6 2 0.667 0.0 1.0 0.0 0.2 1.0 1.0 0.6 1.116
SPONGE-$k$ 5881 0.001 2 0.501 0.0 0.697 0.0 0.001 1.0 5.092 0.003 2.138
$k=4$ LSPCD (avg) 217 0.154 4 0.182 0.12 0.929 0.815 0.143 0.256 23.333 3.586 2.117
LSPCD (std) 13 0.028 0 0.027 0.029 0.004 0.123 0.011 0.009 0.765 0.534 0.408
SCG-MA 176 0.014 4 0.488 0.07 0.914 -0.44 0.172 0.2 25.121 0.349 0.302
SCG-MO 180 0.014 4 0.487 0.07 0.91 -0.431 0.169 0.205 25.252 0.341 0.348
SCG-B 216 0.011 4 0.473 0.052 0.865 0.296 0.076 0.176 12.401 0.142 6.377
SCG-R 450 0.11 4 0.036 0.008 0.92 -0.627 0.033 0.238 8.033 0.886 1.237
KOCG-top-$1$ 26 0.5 4 0.859 0.621 1.0 0.653 0.738 0.112 8.41 4.205
KOCG-top-$r$ 176 0.54 4 0.136 0.041 0.856 -0.246 0.113 0.157 5.034 2.717
BNC-$(k+1)$ 58 0.083 4 0.112 0.0 -0.516 0.0 0.32 0.576 -9.414 -0.784 0.185
BNC-$k$ 5881 0.001 4 0.029 0.0 0.721 0.67 0.001 1.0 5.208 0.004 0.178
SPONGE-$(k+1)$ 71 0.045 4 0.754 0.0 1.0 0.0 0.016 0.443 1.099 0.05 3.364
SPONGE-$k$ 5881 0.001 4 0.75 0.0 0.697 0.0 0.001 1.0 5.092 0.003 2.797
$k=6$ LSPCD (avg) 194 0.114 6 0.251 0.143 0.948 0.646 0.155 0.231 20.031 2.201 2.73
LSPCD (std) 23 0.044 0 0.046 0.053 0.007 0.304 0.019 0.015 1.827 0.582 0.468
SCG-MA 430 0.009 6 0.536 0.021 0.931 -0.355 0.055 0.301 14.568 0.125 0.448
SCG-MO 412 0.009 6 0.571 0.028 0.929 -0.337 0.058 0.3 15.165 0.14 0.477
SCG-B 326 0.017 6 0.313 0.009 0.866 -0.421 0.053 0.222 9.321 0.16 10.509
SCG-R 860 0.038 6 0.038 0.006 0.941 -0.529 0.017 0.367 6.861 0.258 2.125
KOCG-top-$1$ 28 0.429 6 0.867 0.338 1.0 0.197 0.537 0.055 4.071 1.745
KOCG-top-$r$ 430 0.333 6 0.077 0.013 0.88 -0.405 0.043 0.26 3.601 1.2
BNC-$(k+1)$ 224 0.018 6 0.075 0.001 -0.622 0.958 0.033 0.394 -4.239 -0.077 0.286
BNC-$k$ 5881 0.001 6 0.075 0.0 0.722 0.657 0.001 1.0 5.197 0.003 0.194
SPONGE-$(k+1)$ 222 0.016 6 0.622 0.0 1.0 0.0 0.006 0.401 1.252 0.02 1.959
SPONGE-$k$ 5881 0.001 6 0.563 0.0 0.696 -1.0 0.001 1.0 5.085 0.003 2.473
Table 8: Detailed results on the WikiVot dataset from various aspects. LSPCD (avg) and LSPCD (std) respectively indicate the mean and standard deviation across five runs of our method with different seeds.
SIZE BAL K MAC MAO CC+ CC- DENS ISO POL BA-POL RT
$k=2$ LSPCD (avg) 1278 0.224 2 0.038 0.015 0.831 0.673 0.06 0.548 62.322 13.948 2.457
LSPCD (std) 4 0.001 0 0.0 0.0 0.0 0.003 0.0 0.001 0.003 0.056 0.276
SCG-MA 813 0.005 2 0.048 0.079 0.846 0.671 0.104 0.436 71.476 0.353 1.537
SCG-MO 748 0.005 2 0.052 0.082 0.854 0.671 0.113 0.411 71.733 0.385 1.432
SCG-B 414 0.02 2 0.054 0.033 0.756 0.776 0.12 0.221 37.589 0.737 7.501
SCG-R 1100 0.085 2 0.032 0.013 0.83 0.618 0.061 0.444 54.693 4.63 0.781
KOCG-top-$1$ 10 0.5 2 0.905 0.857 1.0 1.0 0.844 0.012 7.6 3.8
KOCG-top-$r$ 813 0.964 2 0.047 0.022 0.427 -0.337 0.066 0.297 2.312 2.229
[…]

### v2

@@ G.8 Results on Real-World Datasets
Table 4: Detailed results for the WoW-EP8 dataset. LSPCD (avg) and LSPCD (std) respectively indicate the mean and standard deviation across five runs of our method with different seeds.
$k$ SIZE IF POL K MAC MAO CC+ CC- DENS ISO TIME (s)
$2$ LSPCD (avg) 586 0.176 223.406 2 0.261 0.098 0.757 0.509 0.511 0.769 0.249
LSPCD (std) 0 0.0 0.0 0 0.0 0.0 0.0 0.0 0.0 0.0 0.034
N2PC ($\gamma=1)$ 527 0.0 236.626 1 0.519 0.0 0.765 0.0 0.588 0.723 26.916
N2PC ($\gamma=1.2)$ 519 0.004 236.543 2 0.763 0.151 0.77 1.0 0.593 0.709 28.426
N2PC ($\gamma=1.5)$ 535 0.053 233.701 2 0.268 0.132 0.767 0.715 0.571 0.724 27.133
N2PC ($\gamma=1.7)$ 552 0.129 227.609 2 0.267 0.109 0.767 0.584 0.542 0.73 33.419
N2PC ($\gamma=2.0)$ 571 0.235 217.398 2 0.27 0.082 0.769 0.48 0.504 0.726 26.603
SCG-MA 527 0.0 236.55 1 0.52 0.0 0.762 0.0 0.59 0.725 1.025
SCG-MO 517 0.0 236.592 1 0.527 0.0 0.769 0.0 0.596 0.708 1.026
SCG-B 583 0.049 200.604 2 0.274 0.094 0.697 -0.369 0.514 0.767 4.657
SCG-R 513 0.03 214.616 2 0.272 0.104 0.761 0.451 0.552 0.654 0.517
KOCG-top-$1$ 16 0.967 13.0 2 0.986 0.889 0.965 0.778 1.0 0.019
KOCG-top-$r$ 527 0.989 12.964 2 0.467 0.11 0.658 -0.598 0.559 0.691
BNC-$(k+1)$ 3 0.792 -0.667 2 0.5 0.0 -1.0 0.0 0.333 0.007 0.99
BNC-$k$ 790 0.005 184.63 2 0.152 0.049 0.628 1.0 0.372 1.0 0.537
SPONGE-$(k+1)$ 375 0.425 87.957 2 0.204 0.095 0.738 0.318 0.343 0.346 0.581
SPONGE-$k$ 790 0.236 191.38 2 0.191 0.093 0.696 0.15 0.372 1.0 0.572
$4$ LSPCD (avg) 590 0.095 218.458 4 0.156 0.082 0.76 0.426 0.506 0.772 0.317
LSPCD (std) 1 0.001 0.142 0 0.003 0.002 0.001 0.016 0.0 0.002 0.044
SCG-MA 599 0.138 205.137 4 0.58 0.13 0.762 -0.32 0.527 0.822 1.198
SCG-MO 568 0.102 213.211 4 0.827 0.141 0.77 -0.349 0.55 0.776 1.176
SCG-B 615 0.0 211.561 1 0.421 0.0 0.693 0.0 0.498 0.822 6.233
SCG-R 503 0.02 214.623 4 0.211 0.056 0.762 0.747 0.564 0.644 1.131
KOCG-top-$1$ 31 0.962 9.054 4 0.962 0.668 0.944 0.339 0.978 0.039
KOCG-top-$r$ 599 0.987 7.393 4 0.436 0.099 0.692 -0.618 0.516 0.806
BNC-$(k+1)$ 8 0.698 -0.25 4 0.5 0.0 -1.0 0.0 0.036 0.003 0.62
BNC-$k$ 790 0.011 185.305 4 0.327 0.03 0.632 1.0 0.372 1.0 0.594
SPONGE-$(k+1)$ 485 0.929 53.823 4 0.448 0.06 0.886 -0.53 0.33 0.433 0.593
SPONGE-$k$ 790 0.869 71.162 4 0.383 0.08 0.803 -0.499 0.372 1.0 0.61
$6$ LSPCD (avg) 591 0.075 217.344 6 0.142 0.071 0.761 0.414 0.505 0.773 0.269
LSPCD (std) 0 0.001 0.142 0 0.006 0.002 0.0 0.009 0.0 0.001 0.015
SCG-MA 598 0.106 207.299 6 0.785 0.113 0.763 -0.339 0.527 0.819 1.226
SCG-MO 591 0.112 205.796 6 0.756 0.171 0.77 -0.321 0.534 0.811 1.378
SCG-B 615 0.0 211.561 1 0.421 0.0 0.693 0.0 0.498 0.822 7.512
SCG-R 744 0.016 201.172 5 0.323 0.178 0.669 0.313 0.41 0.978 1.295
KOCG-top-$1$ 42 0.994 7.905 6 0.992 0.605 0.984 0.303 0.934 0.053
KOCG-top-$r$ 598 0.976 9.109 6 0.472 0.095 0.73 -0.637 0.522 0.812
BNC-$(k+1)$ 10 0.859 -0.2 6 0.5 0.0 -1.0 0.0 0.022 0.002 0.977
BNC-$k$ 790 0.009 185.198 6 0.552 0.016 0.632 1.0 0.372 1.0 0.61
SPONGE-$(k+1)$ 572 0.899 47.762 6 0.537 0.065 0.893 -0.511 0.326 0.536 0.607
SPONGE-$k$ 790 0.878 57.868 6 0.53 0.075 0.834 -0.529 0.372 1.0 0.599
Table 5: Detailed results for the Bitcoin dataset. LSPCD (avg) and LSPCD (std) respectively indicate the mean and standard deviation across five runs of our method with different seeds.
$k$ SIZE IF POL K MAC MAO CC+ CC- DENS ISO TIME (s)
$2$ LSPCD (avg) 155 0.648 29.022 2 0.2 0.143 0.94 0.969 0.199 0.211 2.203
LSPCD (std) 0 0.0 0.013 0 0.0 0.0 0.003 0.001 0.001 0.001 0.161
N2PC ($\gamma=1)$ 134 0.016 29.642 2 0.617 0.195 0.909 0.733 0.246 0.184 11.318
N2PC ($\gamma=1.2)$ 164 0.46 30.146 2 0.238 0.141 0.91 0.964 0.201 0.222 12.26
N2PC ($\gamma=1.5)$ 64 1.0 24.375 2 0.263 0.517 0.941 0.992 0.397 0.133 12.818
N2PC ($\gamma=1.7)$ 70 1.0 23.857 2 0.255 0.452 0.876 0.993 0.364 0.136 13.461
N2PC ($\gamma=2.0)$ 68 1.0 24.147 2 0.258 0.478 0.871 0.993 0.38 0.139 12.834
SCG-MA 179 0.163 28.838 2 0.298 0.068 0.906 0.873 0.179 0.216 0.122
SCG-MO 138 0.032 29.522 2 0.114 0.147 0.91 0.778 0.237 0.184 0.123
SCG-B 40 0.995 21.65 2 0.248 0.87 0.956 1.0 0.56 0.201 2.243
SCG-R 842 0.249 14.24 2 0.022 0.009 0.908 0.812 0.019 0.394 1.024
KOCG-top-$1$ 2 1.0 1.0 2 1.0 1.0 0.0 1.0 1.0 0.004
KOCG-top-$r$ 179 0.992 3.754 2 0.063 0.055 0.266 0.182 0.094 0.165
BNC-$(k+1)$ 50 0.134 -10.76 2 0.058 0.0 -0.516 0.0 0.425 0.581 0.341
BNC-$k$ 5881 0.017 5.268 2 0.059 0.001 0.721 0.694 0.001 1.0 0.202
SPONGE-$(k+1)$ 6 0.792 1.0 2 0.667 0.0 1.0 0.0 0.2 1.0 1.116
SPONGE-$k$ 5881 0.001 5.092 2 0.501 0.0 0.697 0.0 0.001 1.0 2.138
$4$ LSPCD (avg) 217 0.47 23.333 4 0.182 0.12 0.929 0.815 0.143 0.256 2.117
LSPCD (std) 13 0.064 0.765 0 0.027 0.029 0.004 0.123 0.011 0.009 0.408
SCG-MA 176 0.223 25.121 4 0.488 0.07 0.914 -0.44 0.172 0.2 0.302
SCG-MO 180 0.218 25.252 4 0.487 0.07 0.91 -0.431 0.169 0.205 0.348
SCG-B 216 0.233 12.401 4 0.473 0.052 0.865 0.296 0.076 0.176 6.377
SCG-R 450 0.518 8.033 4 0.036 0.008 0.92 -0.627 0.033 0.238 1.237
KOCG-top-$1$ 26 0.905 8.41 4 0.859 0.621 1.0 0.653 0.738 0.112
KOCG-top-$r$ 176 0.931 5.034 4 0.136 0.041 0.856 -0.246 0.113 0.157
BNC-$(k+1)$ 58 0.227 -9.414 4 0.112 0.0 -0.516 0.0 0.32 0.576 0.185
BNC-$k$ 5881 0.01 5.208 4 0.029 0.0 0.721 0.67 0.001 1.0 0.178
SPONGE-$(k+1)$ 71 0.096 1.099 4 0.754 0.0 1.0 0.0 0.016 0.443 3.364
SPONGE-$k$ 5881 0.001 5.092 4 0.75 0.0 0.697 0.0 0.001 1.0 2.797
$6$ LSPCD (avg) 194 0.494 20.031 6 0.251 0.143 0.948 0.646 0.155 0.231 2.73
LSPCD (std) 23 0.153 1.827 0 0.046 0.053 0.007 0.304 0.019 0.015 0.468
SCG-MA 430 0.457 14.568 6 0.536 0.021 0.931 -0.355 0.055 0.301 0.448
SCG-MO 412 0.464 15.165 6 0.571 0.028 0.929 -0.337 0.058 0.3 0.477
SCG-B 326 0.472 9.321 6 0.313 0.009 0.866 -0.421 0.053 0.222 10.509
SCG-R 860 0.407 6.861 6 0.038 0.006 0.941 -0.529 0.017 0.367 2.125
KOCG-top-$1$ 28 0.921 4.071 6 0.867 0.338 1.0 0.197 0.537 0.055
KOCG-top-$r$ 430 0.867 3.601 6 0.077 0.013 0.88 -0.405 0.043 0.26
BNC-$(k+1)$ 224 0.255 -4.239 6 0.075 0.001 -0.622 0.958 0.033 0.394 0.286
BNC-$k$ 5881 0.009 5.197 6 0.075 0.0 0.722 0.657 0.001 1.0 0.194
SPONGE-$(k+1)$ 222 0.147 1.252 6 0.622 0.0 1.0 0.0 0.006 0.401 1.959
SPONGE-$k$ 5881 0.005 5.085 6 0.563 0.0 0.696 -1.0 0.001 1.0 2.473
Table 6: Detailed results for the WikiVot dataset. LSPCD (avg) and LSPCD (std) respectively indicate the mean and standard deviation across five runs of our method with different seeds.
$k$ SIZE IF POL K MAC MAO CC+ CC- DENS ISO TIME (s)
$2$ LSPCD (avg) 1278 0.428 62.322 2 0.038 0.015 0.831 0.673 0.06 0.548 2.457
LSPCD (std) 4 0.002 0.003 0 0.0 0.0 0.0 0.003 0.0 0.001 0.276
N2PC ($\gamma=1)$ 712 0.0 71.635 1 0.11 0.0 0.852 0.0 0.118 0.399 24.313
N2PC ($\gamma=1.2)$ 759 0.006 71.663 2 0.052 0.093 0.847 0.785 0.112 0.418 24.251
N2PC ($\gamma=1.5)$ 760 0.096 70.016 2 0.054 0.029 0.852 0.753 0.109 0.412 22.745
N2PC ($\gamma=1.7)$ 923 0.562 59.142 2 0.056 0.017 0.844 0.685 0.077 0.426 23.51
N2PC ($\gamma=2.0)$ 1190 1.0 40.509 2 0.064 0.012 0.849 0.651 0.042 0.395 39.069
SCG-MA 813 0.008 71.476 2 0.048 0.079 0.846 0.671 0.104 0.436 1.537
SCG-MO 748 0.009 71.733 2 0.052 0.082 0.854 0.671 0.113 0.411 1.432
SCG-B 414 0.037 37.589 2 0.054 0.033 0.756 0.776 0.12 0.221 7.501
SCG-R 1100 0.174 54.693 2 0.032 0.013 0.83 0.618 0.061 0.444 0.781
KOCG-top-$1$ 10 0.717 7.6 2 0.905 0.857 1.0 1.0 0.844 0.012
KOCG-top-$r$ 813 0.999 2.312 2 0.047 0.022 0.427 -0.337 0.066 0.297
BNC-$(k+1)$ 9 0.792 -1.111 2 0.0 0.0 -1.0 0.0 0.139 1.0 0.721
BNC-$k$ 7115 0.003 15.794 2 0.002 0.0 0.558 0.0 0.004 1.0 0.49
SPONGE-$(k+1)$ 10 0.472 1.0 2 0.571 0.0 1.0 0.0 0.111 1.0 1.602
SPONGE-$k$ 7115 0.003 15.794 2 0.057 0.0 0.558 0.0 0.004 1.0 0.977
$4$ LSPCD (avg) 1089 0.519 52.605 4 0.073 0.013 0.856 -0.045 0.072 0.489 4.381
LSPCD (std) 149 0.229 6.003 0 0.026 0.002 0.004 0.427 0.011 0.04 1.69
SCG-MA 1142 0.361 52.945 4 0.081 0.018 0.849 -0.618 0.069 0.506 2.042
SCG-MO 1059 0.374 53.07 4 0.089 0.022 0.858 -0.692 0.073 0.474 1.986
SCG-B 790 0.598 24.782 4 0.091 0.014 0.774 -0.718 0.077 0.342 18.286
SCG-R 1524 0.437 19.524 4 0.031 0.008 0.813 -0.68 0.043 0.549 3.074
KOCG-top-$1$ 33 0.811 4.525 4 0.845 0.086 0.933 -0.609 0.576 0.03
KOCG-top-$r$ 1142 0.99 3.288 4 0.055 0.011 0.719 -0.618 0.059 0.44
BNC-$(k+1)$ 15 0.651 -1.067 4 0.0 0.0 -1.0 0.0 0.076 1.0 0.527
BNC-$k$ 7115 0.001 15.794 4 0.001 0.0 0.558 0.0 0.004 1.0 0.533
SPONGE-$(k+1)$ 12 0.712 1.0 4 0.8 0.0 1.0 0.0 0.091 1.0 2.327
SPONGE-$k$ 7115 0.003 15.794 4 0.156 0.0 0.558 0.0 0.004 1.0 1.522
$6$ LSPCD (avg) 534 0.563 46.179 6 0.143 0.029 0.896 -0.314 0.133 0.287 5.292
LSPCD (std) 46 0.08 2.177 0 0.015 0.002 0.004 0.07 0.012 0.008 0.931
SCG-MA 1355 0.421 45.494 6 0.064 0.023 0.849 -0.647 0.056 0.564 2.24
SCG-MO 1226 0.409 47.013 6 0.073 0.024 0.859 -0.683 0.063 0.526 2.178
SCG-B 941 0.605 23.332 6 0.121 0.018 0.78 -0.735 0.065 0.369 29.072
SCG-R 1501 0.786 10.433 6 0.039 0.008 0.817 -0.734 0.044 0.542 3.475
KOCG-top-$1$ 40 0.963 4.52 6 0.894 0.227 0.981 -0.188 0.564 0.033
KOCG-top-$r$ 1355 0.962 3.132 6 0.051 0.009 0.73 -0.62 0.05 0.506
BNC-$(k+1)$ 13 0.974 -1.077 6 0.0 0.0 -1.0 0.0 0.09 1.0 0.966
BNC-$k$ 7115 0.002 15.794 6 0.001 0.0 0.558 0.0 0.004 1.0 0.546
SPONGE-$(k+1)$ 20 0.859 1.0 6 0.644 0.0 1.0 0.0 0.053 1.0 1.791
SPONGE-$k$ 7115 0.003 15.794 6 0.434 0.0 0.558 0.0 0.004 1.0 1.66
Table 7: Detailed results for the Referendum dataset. LSPCD (avg) and LSPCD (std) respectively indicate the mean and standard deviation across five runs of our method with different seeds.
$k$ SIZE IF POL K MAC MAO CC+ CC- DENS ISO TIME (s)
$2$ LSPCD (avg) 915 0.71 146.109 2 0.279 0.014 1.0 0.114 0.17 0.353 3.376
LSPCD (std) 1 0.002 0.092 0 0.001 0.0 0.0 0.017 0.001 0.001 0.19
N2PC ($\gamma=1)$ 692 0.013 173.604 2 0.542 0.297 1.0 0.573 0.253 0.359 62.32
N2PC ($\gamma=1.2)$ 651 0.02 173.634 2 0.401 0.254 1.0 0.571 0.27 0.343 61.685
N2PC ($\gamma=1.5)$ 918 0.944 130.261 2 0.253 0.011 1.0 0.24 0.149 0.331 70.204
[…]